OpenStudy (anonymous):

Integration! \[\int \frac{x^2 + 4}{x + 2} dx\]

OpenStudy (anonymous):

Ok, I need a u. \(u = x^2 + 4\) \(du = 2x\) Er.. what do I do with the + 2? This could get ugly.

OpenStudy (loser66):

nope

OpenStudy (anonymous):

\(u = x + 2\) maybe?

OpenStudy (loser66):

long division first

OpenStudy (anonymous):

long division works, but I'd complete the square; it's faster

OpenStudy (loser66):

@Euler271 let him do, we gave him the way to do .

OpenStudy (anonymous):

\[= \frac{ (x+2)^2 - 4x }{ x+2 } = x+2 - \frac{ 4x }{ x+2 }\] actually not sure if this is better

OpenStudy (anonymous):

Her, actually. :) And I'm a bit confused at the moment, so... please chime in!

OpenStudy (loser66):

hehehe @Euler271 I found out long division is more beautiful than yours. hehehe

OpenStudy (anonymous):

Do you think it'd help any to break it apart? \[\int \frac{x^2}{x+2} + \frac{4}{x+2} dx\]

OpenStudy (loser66):

nope

OpenStudy (anonymous):

\[= x+2 -\frac{ 4(x+2) - 8 }{ x+2 } = x+2 - 4 +\frac{ 8 }{ x+2 }\]

OpenStudy (loser66):

hey, at the end up, you are back to long division, hehehe @Euler271

OpenStudy (anonymous):

lol XD true its arguable which is faster

OpenStudy (loser66):

yes, it 's fun, right?

OpenStudy (anonymous):

@Euler271 That looks nicer! Thank you! \[\int x - 2 + \frac{8}{x+2} dx\] So that's just \(\frac{x^2}{2} - 2x + \int \frac{8}{x + 2} dx\) (I'll do that one in a sec)

OpenStudy (anonymous):

\[u = x + 2\]\[du = dx\]\[8 \int \frac{1}{u} du = 8\ln |u| + C\] So... \[\frac{x^2}{2} - 2x + \ln |x+2| + C\]

OpenStudy (anonymous):

Oops, with the 8.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you!